5(w+2)-6(w+5)=0
W=

Answer by krutarthas(22)

5(w+2)-6(w+5)=0
⇒5w+10-6w-30=0
⇒5w-6w-30+10=0
⇒-w-20=0
⇒-w = 20
⇒w = -20